# Learning Macroeconomic Policies based on Microfoundations: A Stackelberg Mean Field Game Approach

### v1

@@ Abstract
Effective macroeconomic policies play a crucial role in promoting economic growth and social stability. This paper models the optimal macroeconomic policy problem based on the Stackelberg Mean Field Game (SMFG), where the government acts as the leader in policy-making, and large-scale households dynamically respond as followers. This modeling method captures the asymmetric dynamic game between the government and large-scale households, and interpretably evaluates the effects of macroeconomic policies based on microfoundations, which is difficult for existing methods to achieve. We also propose a solution for SMFGs, incorporating pre-training on real data and a model-free Stackelberg mean-field reinforcement learning (SMFRL) algorithm, which operates independently of prior environmental knowledge and transitions. Our experimental results showcase the superiority of the SMFG method over other economic policies in terms of performance, efficiency-equity tradeoff, and SMFG assumption analysis. This paper significantly contributes to the domain of AI for economics by providing a powerful tool for modeling and solving optimal macroeconomic policies.
Macroeconomic policies; Stackelberg mean field game; reinforcement learning; multi-agent; social good
@@ 1 Introduction
The formulation of macroeconomic policies is crucial for the sustained development of an economy (Schneider & Frey, 1988; Persson & Tabellini, 1999). Governments can adjust economic production, wealth distribution, social stability, and welfare through economic policies such as interest rates, taxation, and fiscal spending. Thus, how to effectively model and solve optimal macroeconomic policies, and simulate the effect of policy implementation is a very important issue. Existing research methods in economic policy primarily rely on simulations based on macroeconomic models like the Dynamic Stochastic General Equilibrium (DSGE) model (An & Schorfheide, 2007) and the Solow growth model (Brock & Taylor, 2010), or on empirical analysis (Robnik-Šikonja & Kononenko, 2003) assessing policy impacts through posterior analysis of historical data. However, these methods have limitations: (1) Difficulty in modeling dynamic interactions between the government and large-scale households, including feedback from households to policy and governmental adjustments in response to this feedback. (2) Challenge in solving the behavioral strategies of large-scale micro-agents. (3) The real economic and social world is highly complex, making it challenging for macroeconomic models to satisfy assumptions such as state transition and complete information.
Addressing these challenges, this paper employs the Stackelberg Mean Field Game (SMFG) to model optimal macroeconomic policy problems. In this framework, the government, acting as the leader, initially sets policies, while a vast number of households, as followers, respond to these policies. However, considering the large scale of the followers, it is impractical for the leader to consider each follower’s dynamic response. Consequently, followers’ action-state distribution, analogous to the overall market information in economics, is introduced as a bridge between macro-level policy and micro-level individual decisions. The optimal decisions of the leader and followers depend on the followers’ overall distribution rather than on a specific agent. This assumption is crucial for modeling optimal macroeconomic policies as SMFGs and is easily comprehensible in the real world. For example, the government lowers property taxes to stimulate a sluggish housing market. This policy alters the market’s overall state, including housing prices and market demand. Households should consider both the government’s tax policy and the overall market information when making their purchasing decisions.
Furthermore, this paper proposes a solution method for the optimal macroeconomic policy issues based on SMFG: Firstly, behavior cloning of households (followers) based on real data serves as the initialization for the policy network. This pre-training enhances the stability and performance of training and prevents ineffective solutions. Subsequently, a model-free algorithm, Stackelberg Mean-Field Reinforcement Learning (SMFRL), is introduced to solve SMFGs. This method does not rely on prior knowledge of the environment or transition information, offering advantages in solving complex economic issues. We conduct experiments on TaxAI (Mi et al., 2023), an environment for studying optimal tax policies, first to validate the performance of SMFRL against other model-free algorithms such as independent learning, MADDPG, and mean field MARL in solving SMFGs. Secondly, the paper compares the SMFG method with AI Economist, the Saez tax, the 2022 U.S. Federal tax, and free market policy, demonstrating the superiority of SMFG solutions and exploring the efficiency-equity tradeoff of these economic policies. Lastly, our tests on SMFG modeling assumptions reveal that our approach remains effective even if some households’ optimal decisions do not depend on market information. In summary, the contributions of this paper are threefold:
1. This paper is the first to model optimal macroeconomic policy problems using Stackelberg mean field games, linking macro-policies (the leader) with microfoundations (large-scale followers) through the followers’ overall distribution or average actions. It overcomes limitations such as dynamic interactions and feedback, and the dimensionality curse of large-scale households.
2. This paper employs real data for pre-training and proposes a model-free Stackelberg mean field reinforcement learning algorithm to solve SMFGs, with experimental validation of its effectiveness.
3. The paper empirically showcases the superiority of modeling and solving optimal macroeconomic policy problems by SMFGs. We also explore the efficiency-equity tradeoff in various economic policies and SMFG assumptions.
@@
Our code is shown in an anonymous GitHub repository  [^1] https://anonymous.4open.science/r/SMFG_macro_7740.
@@ 2.1 Economic Methods
In the field of economics, research on economic policy typically relies on theoretical model studies, empirical analysis, and econometric methods. Classic theoretical models, such as the IS-LM model (Hicks, 1980; Gali, 1992), are applicable for analyzing short-term policy effects but overlook long-term factors and price fluctuations. The aggregate demand-aggregate supply (AD-AS) model (Dutt, 2006; Lee et al., 1997) focuses on the relationship between aggregate demand and supply, integrating short-term and long-term factors, but simplifies macroeconomic dynamics. The Solow growth model (Brock & Taylor, 2010) provides a framework for economic growth but does not consider market imperfections and externalities. New Keynesian models (Blanchard & Galí, 2007; Gabaix, 2020) emphasize the stickiness of prices and wages, suitable for explaining economic fluctuations and policy interventions, but are based on strong assumptions. DSGE models (An & Schorfheide, 2007; Smets & Wouters, 2007), built on micro-foundations for macroeconomic forecasting, offer a consistent framework for policy analysis but are limited in handling nonlinearities and market imperfections. Empirical analysis (Ramesh et al., 2010; Vedung, 2017) primarily evaluates policy effects through historical data but faces limitations in data timeliness and accuracy. Econometric methods (Johnston & DiNardo, 1963; Davidson et al., 2004), such as regression and time series analysis, can quantify policy effects but are constrained by model settings and data quality.
@@ 2.2 AI for Economics
Artificial intelligence (AI), such as reinforcement learning, may offer new perspectives for solving complex economic problems (Tilbury, 2022). In terms of macroeconomics, AI Economist (Zheng et al., ) employs curriculum learning on tax policy design, Trott et al. studies the collaboration between central and local governments under COVID-19, as well as monetary policy (Hinterlang & Tänzer, ; Chen et al., 2023), international trade (Sch, 2021), market pricing (Danassis et al., 2023). However, these studies consider the government as a single agent, neglecting the dynamic response of the households to policies. Koster et al. studies democratic AI by human voting models, Yaman et al. examines the impact of social sanction rules on labor division. However, these works are based on simplified settings and remain a gap in real-world economic policy. In terms of microeconomics, Shi (2021); Rui & Shi (2022); Atashbar & Aruhan Shi (2023) study the optimal saving and consumption problems of micro-agents, and other researchers explore rational expectation equilibrium (Kuriksha, 2021; Hill et al., ), multiple equilibria under real-business-cycle (Curry et al., 2022), the emergence of barter behavior (Johanson et al., 2022), optimal asset allocation and savings strategies (Ozhamaratli & Barucca, 2022). In summary, research on macroeconomic problems still lacks a bridge framework of macro policy and microfoundations, which has drawn our attention to the Stackelberg mean field game.
@@ 2.3 Stackelberg Mean Field Game
The Stackelberg Mean Field Game models one leader and a large number of homogeneous followers, and current research can be divided into model-based methods and model-free methods. Model-based approaches typically involve complex theoretical analysis and precise mathematical modeling. For instance, Guo et al. (2022) reconceptualizes SMFGs as minimax optimization problems, while Dayanikli & Lauriere (2023); Bergault et al. (2023); Bensoussan et al. (2015); Fu & Horst (2020) initially resolve the followers’ mean field equilibrium using forward-backward stochastic functional differential equations, subsequently deriving the leader’s optimal control. Huang & Yang (2020); Bensoussan et al. (2015) explore linear-quadratic Stackelberg games. These methods provide a theoretically robust foundation but may require simplified assumptions for practical applications, particularly in complex real-world scenarios. Notably, Aurell et al. (2022) extends the mean field game model to capture epidemic dynamics, demonstrating potential applications in public health management. In contrast, model-free methods offer a way to handle complex situations that traditional numerical methods struggle with, especially in data-rich environments. Campbell et al. (2021) employs a modified deep backward stochastic differential equation (BSDE) method for solving the followers’ equilibrium, yet the leader’s choices are limited. In conclusion, we plan to use SMFGs to model macroeconomic policy issues and propose a model-free algorithm to solve this complex task.
Figure 1: The mapping from optimal macroeconomic policy problems (left) to the Stackelberg Mean Field Games (right). The left represents the reality where the macro-level government implements fiscal or monetary policies to optimize GDP, welfare, or equity, and the micro-level consists of large-scale households engaging in economic activities and forming the market. In SMFGs, the government corresponds to the leader, households to the followers, and market information to the followers’ state-action distribution.
3 Macroeconomic Policy Formulation
3.1 Economic Modeling
In reality, an economy comprises at least one government and large-scale households. The diverse economic activities of households, such as production, trade, and consumption, aggregate to form the market. The government can maintain economic growth and social stability (Figure 1 left) via economic policies. The following is a mathematical model for optimizing tax policy:
Households
At timestep $t$, a household holds assets $a_{t}$. Its income $i_{t}$ consists of labor income $W_{t}e_{t}h_{t}$, and capital income $r_{t}a_{t}$. Here, $W_{t}$ and $r_{t}$ are the wage and interest rates determined by the labor and capital markets, respectively, while $e_{t}$ indicates the private education level. Households are obligated to pay taxes as per the government’s tax policy $T_{t}(i_{t},a_{t})$. Each household decides on its consumption $c_{t}\in R^{+}$, working hours $h_{t}\in R^{+}$, and next savings $a_{t+1}\in R^{+}$ to maximize long-term utility until $T_{max}$, subject to the budget constraint:
$\max_{a_{t+1},h_{t}}$ $\quad\mathbb{E}_{0}\sum_{t=0}^{T_{max}}u(c_{t},h_{t})$

### v2

# Learning Macroeconomic Policies based on Microfoundations: A Dynamic Stackelberg Mean Field Game Approach

@@ Abstract
The Lucas critique emphasizes the importance of considering the impact of policy changes on the expectations of micro-level agents in macroeconomic policymaking. However, the inherently self-interested nature of large-scale micro-agents, who pursue long-term benefits, complicates the formulation of optimal macroeconomic policies. This paper proposes a novel general framework named Dynamic Stackelberg Mean Field Games (Dynamic SMFG) to model such policymaking within sequential decision-making processes, with the government as the leader and households as dynamic followers. Dynamic SMFGs capture the dynamic interactions among large-scale households and their response to macroeconomic policy changes. To solve dynamic SMFGs, we propose the Stackelberg Mean Field Reinforcement Learning (SMFRL) algorithm, which leverages the population distribution of followers to represent high-dimensional joint state and action spaces. In experiments, our method surpasses macroeconomic policies in the real world, existing AI-based and economic methods. It allows the leader to approach the social optimum with the highest performance, while large-scale followers converge toward their best response to the leader’s policy. Besides, we demonstrate that our approach retains effectiveness even when some households do not adopt the SMFG policy. In summary, this paper contributes to the field of AI for economics by offering an effective tool for modeling and solving macroeconomic policy-making issues.
@@ 1 Introduction
Macroeconomic policymaking is fundamental to the sustained development of an economy [52, 46]. Governments influence economic production, wealth distribution, social stability, and welfare through economic policies such as interest rates, taxation, and fiscal spending [49]. Therefore, accurately modeling and solving optimal macroeconomic policies and simulating the effects of their implementation are critical issues. According to the Lucas critique [39], predicting the impact of macroeconomic policies effectively requires modeling individual micro-level behavior—termed "microfoundations"—to forecast personal responses to policy changes and aggregate these into macroeconomic effects. Nonetheless, several challenges persist: dynamic interactions among micro-agents, between micro-agents and macro-authorities, the inherently selfish behavior of agents focused on optimizing long-term returns, and the curse of dimensionality due to the large-scale nature of agents such as households and firms.
This paper introduces a novel framework termed Dynamic Stackelberg Mean Field Game (Dynamic SMFG) designed to address complex macroeconomic policy-making problems. Within this dynamic setting, the government, serving as the leader, adjusts policies iteratively, while a large population of households, acting as followers, respond accordingly. This process involves sequential decision-making, with both the leader and the followers aiming to optimize their cumulative long-term benefits. Given the leader’s policy, the SMFG reduces to a mean field game, which is widely used for modeling interactions among large-scale homogenous agents [33]. In Dynamic SMFGs, followers aim to maximize their long-term benefits by providing the best response to the leader’s policy, while the leader’s goal is to find an optimal policy that effectively guides the massive followers toward achieving comprehensive macroeconomic objectives. Existing approaches often simplify Dynamic SMFGs to single-step decisions [45] or assume myopic followers [63], which do not adequately capture the complexities of macroeconomic policy-making. Therefore, we propose a data-driven reinforcement learning algorithm, Stackelberg Mean-Field Reinforcement Learning (SMFRL), to solve Dynamic SMFGs without requiring transition information or prior knowledge. The SMFRL algorithm utilizes a mean field approximation [28] to reduce the high-dimensional joint state and actions of followers to a manageable population distribution, allowing for the effective learning of a central critic Q function based on the overall followers’ distribution.
Experiments conducted on TaxAI [41], an environment designed for studying optimal tax policies, first validate the performance of SMFRL in solving Dynamic SMFGs. The results indicate that SMFRL enables followers to closely achieve their best response while allowing the leader to approach the social optimum with the highest performance compared with baselines. Furthermore, the paper compares the Dynamic SMFG method with macroeconomic policies such as AI Economist [62], the Saez tax [50], the 2022 U.S. Federal tax, and free market policy [5], demonstrating the superiority of the Dynamic SMFG framework in macroeconomic policymaking. Considering real-world applications, our experiments show that even if a portion of households do not adopt the follower policy learned by the SMFG method, it does not adversely affect the utility, wealth, and income of households that do adopt the SMFG policy, which are in fact higher. However, the GDP of the economy may decrease with an increase in the number of households not supporting the SMFG policy, as they typically contribute less labor. In summary, the contributions of this paper are threefold:
1. It innovatively proposes a general framework based on Dynamic Stackelberg Mean Field Games to model macroeconomic policy-making, considering key issues such as microfoundations, dynamic interactions and feedback, rational expectations, and randomness.
2. It introduces a data-driven Stackelberg Mean Field Reinforcement Learning algorithm to solve dynamic SMFGs without environmental transition information and prior knowledge, validated experimentally for its effectiveness.
3. It empirically demonstrates the superiority of the SMFG method in modeling and solving optimal macroeconomic policy problems and discusses its robustness in real-world scenarios.
@@
The source code for the SMFG method is available in an anonymous GitHub repository  [^1] https://anonymous.4open.science/r/SMFG_macro_7740.
@@ 2.1 Economic Methods
In the field of economics, research on economic policy typically relies on theoretical model studies, empirical analysis, and econometric methods. Classic theoretical models, such as the IS-LM model [24, 22], are applicable for analyzing short-term policy effects but overlook long-term factors and price fluctuations. The aggregate demand-aggregate supply (AD-AS) model [19, 34] focuses on the relationship between aggregate demand and supply, integrating short-term and long-term factors, but simplifies macroeconomic dynamics. The Solow growth model [10] provides a framework for economic growth but does not consider market imperfections and externalities. New Keynesian models [9, 21] emphasize the stickiness of prices and wages, suitable for explaining economic fluctuations and policy interventions, but are based on strong assumptions. DSGE models [1, 55], built on micro-foundations for macroeconomic forecasting, offer a consistent framework for policy analysis but are limited in handling nonlinearities and market imperfections. Empirical analysis [47, 59] primarily evaluates policy effects through historical data but faces limitations in data timeliness and accuracy. Econometric methods [30, 15], such as regression and time series analysis, can quantify policy effects but are constrained by model settings and data quality.
@@ 2.2 AI for Economics
Artificial intelligence (AI), such as reinforcement learning, may offer new perspectives for solving complex economic problems [56]. In terms of macroeconomics, AI Economist [62] employs curriculum learning on tax policy design, [58] studies the collaboration between central and local governments under COVID-19, as well as monetary policy [26, 12], international trade [51], market pricing [14]. However, these studies consider the government as a single agent, neglecting the dynamic response of the households to policies. [31] studies democratic AI by human voting models, [60] examines the impact of social sanction rules on labor division. However, these works are based on simplified settings and remain a gap in real-world economic policy. In terms of microeconomics, [53, 48, 3] study the optimal saving and consumption problems of micro-agents, and other researchers explore rational expectation equilibrium [32, 25], multiple equilibria under real-business-cycle [13], the emergence of barter behavior [29], optimal asset allocation and savings strategies [44]. In summary, research on macroeconomic problems still lacks a bridge framework of macro policy and microfoundations, which has drawn our attention to the Stackelberg mean field game.
Figure 1: The mapping from optimal macroeconomic policy problems (left) to the Dynamic Stackelberg Mean Field Games (right). The left represents the reality where the macro-level government implements fiscal or monetary policies to optimize GDP, welfare, or equity, and the micro-level consists of large-scale households engaging in economic activities. In dynamic SMFGs, the government acts as the leader, and households as followers, respond to the leader’s policy.
@@ 2.3 Stackelberg Mean Field Game
Stackelberg Mean Field Games (SMFG) model interactions between the leader and numerous homogeneous followers, employing both model-based and model-free approaches. Model-based methods, such as those presented in [23], treat SMFGs as minimax optimization problems, streamlining the leader’s strategy to a single optimal action rather than a comprehensive policy. Studies by [7, 20, 16, 8] utilize forward-backward stochastic functional differential equations to determine the followers’ mean field equilibrium before computing the leader’s optimal strategy. This includes addressing terminal state constraints [20] and simplifying scenarios via linear-quadratic formulations [6, 37, 43, 18, 27]. Such methods, though theoretically robust, often rely on simplifying real-world complexities and require extensive knowledge of environmental transitions. Extending SMFG to model epidemic dynamics [4] showcases its broader applicability in public health management. Conversely, model-free approaches leverage complex, data-rich environments without the need for prior environmental knowledge. [11] employs a modified deep backward stochastic differential equation framework to establish followers’ equilibrium, despite constraints on leader decisions. [45] investigates applications in privacy protection through a single-step decision model. [42] delves into locally optimal defensive strategies given predefined attacker trajectories. [35] learns followers’ transition functions from empirical data, and facilitates SMFG solutions through the empirical Fokker-Planck equation.
However, in real-world settings, the sequential decision-making between a leader and numerous followers to optimize long-term gains is common but inadequately studied. This paper proposes a data-driven reinforcement learning algorithm for solving the aforementioned problem, termed the Dynamic Stackelberg Mean Field Game, without reliance on environmental transition information or prior knowledge.
3 Stackelberg Mean Field Game Modeling
3.1 Macroeconomic Policy-Making
Macroeconomic policy-making involves formulating effective policies to regulate and guide an economy toward achieving sustainable growth, low unemployment, and stable prices. In addressing these challenges, several key elements must be considered: ① Microfoundations: According to Lucas critique, basing macroeconomic models on microeconomic behaviors is crucial. This ensures that the actions of individual agents, such as households and firms, aggregate to produce macroeconomic outcomes. ② Dynamic Interactions: Extensive interactions occur both between the macro-agent (the government) and a multitude of micro-agents (households), and within groups of households. These interactions can lead to complex economic dynamics and endogenous shocks. ③ Rational Expectations: Agents are typically modeled as rational and self-interested, making decisions based on available information and their expectations about future policies and economic conditions. ④ Stochastic Elements: Incorporating randomness and shocks, such as technological changes, policy shifts, and external economic shocks, is essential for understanding variability in economic outcomes. ⑤ Feedback Mechanisms: Feedback loops exist within the economy where macro-policy impacts micro-agents, and the micro-agents’ behaviors, in turn, influence policy decisions.
Given the aforementioned key points, Dynamic Stackelberg Mean Field Games (Dynamic SMFGs) are ideal for macroeconomic policy-making due to:
1. Leader-Follower Dynamics (②⑤): The government, acting as the leader, sets policies first. Households, as followers, adjust their behaviors based on these policies, which captures the strategic interactions between policy-setters and economic agents.
2. Mean Field Interaction (①⑤): The collective behavior of numerous micro-agents is efficiently modeled as a "mean field", which simplifies the analysis of large-scale interactions and their overall impact on the macroeconomy.
3. Dynamic Strategy (②③④): The dynamic nature of these games reflects the evolving economic conditions and policy-making, enabling the model to adapt to new information and economic shocks.
3.2 Dynamic Stackelberg Mean Field Games
Following this, we consider a dynamic SMFG over discrete periods $t\in\{0,\ldots,T\}$ for macroeconomic policy-making issues: At each time $t$, the leader first selects an action $a_{t}^{l}\in\mathcal{A}^{l}$ based on the leader’s state $s_{t}^{l}\in\mathcal{S}^{l}$ and policy $\pi^{l}:\mathcal{S}^{l}\rightarrow\mathcal{A}^{l}$. Subsequently, followers simultaneously determine their actions based on the leader’s action $a^{l}_{t}$ and their private states $s^{f}_{t}\in\mathcal{S}^{f}$. The representative follower’s action $a_{t}^{f}\in\mathcal{A}^{f}$ is derived from a shared policy $\pi^{f}:\mathcal{S}^{f}\times\mathcal{A}^{l}\rightarrow\mathcal{A}^{f}$. The sequences $\{\pi_{t}^{l}\}_{t=0}^{T}$ and $\{\pi_{t}^{f}\}_{t=0}^{T}$ are denoted as $\pi^{l}$ and $\pi^{f}$, respectively.
Followers’ Mean Field Game
Given the leader’s policy $\pi^{l}$, the dynamic SMFG is simplified into a mean field game for the followers. At time $t\in\{0,...,T-1\}$, given the leader’s action $a^{l}_{t}$, the representative follower’s state $s^{f}_{t}$ and action $a^{f}_{t}$, and the population state-action distribution of all followers denoted as $L^{f}_{t}\in\mathcal{P}(\mathcal{S}^{f}\times\mathcal{A}^{f})$, the representative follower receives a reward $r^{f}(s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$ and move to the next state $s_{t+1}^{f}\sim\mathbb{P}(\cdot|s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$. Each follower aims to find the optimal policy $\pi^{f}$ that maximizes his cumulative reward over the time horizon:
$J^{f}(\pi^{l},\pi^{f},L^{f})=\mathbb{E}^{f}\left[\sum_{t=0}^{T}r^{f}(s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})\right]$ (1)
where $\mathbb{E}^{f}$ denotes the expectation with respect to initial state $s_{0}^{f}\in\mu_{0}^{f}$, $(s_{t+1}^{l},s_{t+1}^{f})\sim P(\cdot\mid s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$ for $t\in\{0,...,T-1\}$, and $a^{l}_{t}\sim\pi^{l}(\cdot\mid s_{t}^{l}),a^{f}_{t}\sim\pi^{f}(\cdot\mid s_{t}^{f},a^{l}_{t})$ for $t\in\{0,...,T\}$.
Definition 3.1.
[Followers’ Best Response for Leader’s Policy] Given any leader’s policy $\pi^{l}\in\Pi^{l}$ and the followers’ state-action distribution $L^{f}\in\mathcal{P}(\mathcal{S}^{f}\times\mathcal{A}^{f})$, the followers’ best response policy ${\pi^{f}}^{*}(\pi^{l},L^{f})$ is defined as
${\pi^{f}}^{*}(\pi^{l},L^{f})\in\arg\max_{\pi^{\prime}}J^{f}(\pi^{l},\pi^{\prime},L^{f}).$
Leader’s Game
At time $t\in\{0,...,T-1\}$, given the leader’s state $s^{l}_{t}$, action $a^{l}_{t}$, and the followers’ state-action distribution $L_{t}^{f}$, the leader will receive a reward $r^{l}(s_{t}^{l},a_{t}^{l},L^{f}_{t})$, and move to a new state $s_{t+1}^{l}$ according to the transition probabilities $\mathbb{P}(\cdot|s_{t}^{l},a_{t}^{l},L^{f}_{t})$. The leader aims to find the optimal policy $\pi^{l}$ to maximize the total expected reward:
$J^{l}(\pi^{l},\pi^{f},L^{f})=\mathbb{E}^{l}\left[\sum_{t=1}^{T}r^{l}(s_{t}^{l},a_{t}^{l},L^{f}_{t})\right],$ (2)
where $\mathbb{E}^{l}$ denotes the expectation under this probability distribution for the leader when $s_{0}^{l}\sim\mu_{0}^{l},s_{t+1}^{l}\sim\mathbb{P}(\cdot|s_{t}^{l},a_{t}^{l},L^{f}_{t})$ for $t\in\{0,...,T-1\}$, and $a_{t}^{l}\sim\pi_{t}^{l}(\cdot\mid s_{t}^{l})$ for $t\in\{0,...,T\}$.
Definition 3.2 (Leader’s Optimal Policy in Dynamic Stackelberg Mean Field Games).
Considering the followers’ best response $(\pi^{f},L^{f})$ for the leader’s policy, which satisfies Definition 3.1, the leader’s optimal policy ${\pi^{l}}^{*}$ in dynamic Stackelberg mean field games is defined as follows:
${\pi^{l}}^{*}$ $\in\mathop{\arg\max}_{{\pi^{l}}^{\prime}}J^{l}({\pi^{l}}^{\prime},\pi^{f},L^{f}),$